# PhysBrain: Human Egocentric Data as a Bridge from Vision Language Models to Physical Intelligence

### v1

@@ Abstract
Robotic generalization relies on physical intelligence: the ability to reason about state changes, contact-rich interactions, and long-horizon planning under egocentric perception and action. However, most VLMs are trained primarily on third-person data, creating a fundamental viewpoint mismatch for humanoid robots. Scaling robot egocentric data collection remains impractical due to high cost and limited diversity, whereas large-scale human egocentric videos offer a scalable alternative that naturally capture rich interaction context and causal structure. The key challenge is to convert raw egocentric videos into structured and reliable embodiment training supervision. Accordingly, we propose an Egocentric2Embodiment translation pipeline that transforms first-person videos into multi-level, schema-driven VQA supervision with enforced evidence grounding and temporal consistency, enabling the construction of the Egocentric2Embodiment dataset (E2E-3M) at scale. An egocentric-aware embodied brain, termed PhysBrain, is obtained by training on the E2E-3M dataset. PhysBrain exhibits substantially improved egocentric understanding, particularly for planning on EgoThink. It provides an egocentric-aware initialization that enables more sample-efficient VLA fine-tuning and higher SimplerEnv success rates (53.9%), demonstrating effective transfer from human egocentric supervision to downstream robot control.
@@ 1 Introduction
Figure 1: Human egocentric supervision improves first-person embodied brains and transfers to control. Left: EgoThink radar plot comparing egocentric VLM performance across six dimensions (Activity, Forecast, Localization, Object, Planning, Reasoning) for representative baselines. Right Top: "Phys" means that the VLM was supervised fine-tuning on our annotated first-person (egocentric) data (described in Sec. 3.1), both VST-7B and Qwen2.5-VL-7B achieve significantly better EgoThink performance, with the most pronounced gains on Planning. Right Bottom: when used as the VLM backbone in a standard VLA fine-tuning pipeline, the same Phys-enhanced backbones yield substantially higher SimplerEnv success rates, indicating that better egocentric planning and interaction reasoning translate to improved downstream manipulation.
@@
Vision-Language-Action (VLA) systems rely on a reliable embodied brain that integrates scenario understanding and action generation. Recent multimodal systems (Hurst et al., 2024a; Bai et al., 2025a) show rapid gains in visual perception, spatial and video reasoning, and long context understanding. These advances provide rich open vocabulary recognition and semantic inference capabilities that can be transferred to action prediction, thereby enabling modern VLAs (Zitkovich et al., 2023; Kim et al., 2024; Bjorck et al., 2025; Black et al., 2024, 2025) to achieve strong performance across diverse manipulation tasks. These developments highlight that strong VLA performance is driven by an embodied brain that grounds executable planning and interaction decisions in the agent’s own perceptual stream.
@@
For future humanoid robots, this perceptual stream is expected to be predominantly first-person, since perception, planning, and action feasibility are fundamentally grounded in the agent’s own body and workspace (Grauman et al., 2022). This places stringent demands on multimodal models operating under egocentric settings. However, empirical results on egocentric benchmarks (Lin et al., 2022; Pramanick et al., 2023; Chen et al., 2024; Patel et al., 2025; Li et al., 2025a) indicate that current multimodal models still struggle with long-horizon understanding, planning, and reliability under egocentric videos. These deficits stem from challenges intrinsic to egocentric perception, including rapid viewpoint changes, frequent hand–object occlusions, the absence of the actor’s full body, and the need for cross-frame inference of contact and object state (He et al., 2025). Consequently, current performance bottlenecks are more likely due to insufficient egocentric embodied cognition, state tracking, and planning supervision, rather than limitations in model scale or single-frame recognition.
@@
These limitations raise a fundamental scalability question: whether advancing VLA in head-mounted egocentric settings necessarily depends on extensive robot data, including robot egocentric supervision. Acquiring large-scale and diverse robot manipulation data is widely acknowledged to be costly and difficult to scale, due to substantial hardware, labor, and safety constraints (Khazatsky et al., 2024). Even imitation learning relies on expensive human demonstrations, while existing large-scale robot data pipelines often require long collection cycles or sustained multi-institution collaboration (Brohan et al., 2022; Zitkovich et al., 2023; O’Neill et al., 2024). As a result, learning and aligning embodied brains primarily through such robot data fundamentally constrains the scalability and coverage of egocentric VLA systems.
@@
In contrast to costly and hard-to-scale robot data, human first-person videos provide a naturally scalable source of egocentric supervision, covering diverse everyday behaviors and environments. This data modality offers observations closely aligned with real interaction distributions for learning embodied brains. Large-scale datasets such as Ego4D (Grauman et al., 2022), BuildAI (BuildAI, 2025), and EgoDex (Hoque et al., 2025) demonstrate that egocentric videos can capture long-horizon activities, human–object interactions, and fine-grained manipulation dynamics at scale. An open question is how to leverage the latent planning structure and hand–object interaction regularities in human egocentric videos as supervision to strengthen egocentric embodied brains without robot data, thereby improving the sample efficiency and generalization of VLA systems.
@@
Motivated by this observation, we develop a scalable annotation and instruction pipeline that transforms human egocentric videos into structured, multi-level first-person VQA supervision for embodied brain learning. Each VQA instance encodes complementary information across multiple levels, including planning decompositions, key states, interaction constraints, and temporal relations, providing supervision beyond static visual recognition. To directly assess the effectiveness of this supervision, we conduct a controlled evaluation using EgoDex-derived VQA data alone, as shown in Fig.1. Embodied brains trained on top of different VLM backbones consistently outperform their corresponding base models when evaluated as embodied brains. Under this setting, the resulting models enable efficient few-shot adaptation on first-person VLA tasks and achieve performance comparable to, or exceeding, VLA systems trained with large-scale robot data, despite the absence of any robot-data pretraining.
Building on this evidence, we train PhysBrain by scaling the supervision to a mixture of Ego4D, BuildAI, and EgoDex, together with general-purpose vision–language data, to further strengthen egocentric planning and interaction reasoning while preserving general vision–language capability. This direction is complementary rather than a replacement for robot data: robot egocentric supervision remains critical for physical grounding and can further raise the performance ceiling when combined with our approach.
In summary, our contributions are as follows:
@@
  • We introduce a scalable annotation and instruction pipeline, called Egocentric2Embodiment Translation Pipeline, which converts large-scale human egocentric videos from multiple scenarios into multi-level embodied supervision.
@@
  • We find that human egocentric data is complementary to robot data and is significantly more scalable, offering a promising basis for studying future scaling laws in first-person VLA.
@@ 2 Related Work
2.1 First Person Vision Language Model
Vison Language Models (VLMs) that excel on third-person content often degrade when the input shifts to egocentric imagery and video. Multiple lines of evidence point to a persistent viewpoint domain gap and to missing egocentric cues such as hand manipulation, egomotion, and partial observability (He et al., 2025). EgoVLP (Lin et al., 2022) were among the first to document that third-person pretraining transfers poorly and that explicitly egocentric objectives are needed for first-person retrieval, recognition, and temporal grounding. EgoVLPv2 (Pramanick et al., 2023) further reports that fusing first-person video and language during pretraining is important for egocentric tasks. Beyond these early works, recent evaluations arrive at the same conclusion. EgoPlan-Bench (Chen et al., 2024) shows that mainstream multimodal models struggle with egocentric planning even when the scenes are household and the instructions are simple, and it analyzes typical failure modes such as viewpoint confusion and missing contact reasoning. Studies on QaEgo4D (Barmann and Waibel, 2022) and QaEgo4Dv2 (Patel et al., 2025) find that both proprietary and open source VLMs lag on long-horizon egocentric reasoning. EgoM2P (Li et al., 2025a) also emphasizes the structural gap between third-person and first-person streams and argues for egocentric priors during pretraining.
@@ 2.2 Vision Language Action
Vision-Language-Action (VLA) models (Brohan et al., 2023; Zitkovich et al., 2023; Team et al., 2024) represent a recent paradigm shift in robotic manipulation by unifying language understanding, visual perception, and motor control within a single end-to-end framework. Building upon large-scale vision-language models, VLAs directly map high-dimensional visual observations and natural language instructions to low-level robot actions, enabling intuitive human-robot interaction and task execution. Early works such as RT-1 (Brohan et al., 2023) and RT-2 (Zitkovich et al., 2023) demonstrate that scaling robot data and leveraging pretrained vision-language representations significantly improve manipulation performance across diverse tasks. Building upon these foundations, OpenVLA (Kim et al., 2024), $\pi_{0}$ (Black et al., 2024; Pertsch et al., 2025; Black et al., 2025), and GR00T-N1 (Bjorck et al., 2025) further advance VLA capabilities through large-scale cross-embodiment and multi-task pretraining, demonstrating superior generalization and action prediction performance. Several works (Zhou et al., 2025; Yang et al., 2025c; Fang et al., 2025; Mazzaglia et al., 2025) attempt to address the catastrophic forgetting of language capabilities during VLA training, while others (Zawalski et al., 2025; Sun et al., 2024; Lin et al., 2025; Huang et al., 2025; Lee et al., 2025; Yuan et al., 2025) explore incorporating chain-of-thought reasoning into the VLA inference process. To pursue better generalization, several works (Shen et al., 2025; Cen et al., 2025; Liang et al., 2025; Jia et al., 2025) attempt to incorporate video generation models or world models into VLA action prediction, while others (Li et al., 2025b; Yu et al., 2025; Chen et al., 2025a, b) explore applying reinforcement learning to train VLA models. However, the aforementioned works primarily rely on robot-specific data for VLA training. Due to the high cost of robot data collection in practical scenarios, high-quality robot demonstration data remains extremely scarce. In contrast to these approaches, our work explores leveraging human egocentric data for model training, with the hypothesis that large-scale human demonstration data can effectively elicit generalization capabilities in VLA models.
@@ 2.3 Learning VLAs from Human demonstration
Robot data acquisition is hard to scale due to the stringent robot–operator configuration and reliance on expert tele-operation. Egocentric VLA trained on the egocentric human demonstrations offers a more scalable path, with strong potential to advance perception–action learning and real-world executability. EgoVLA (Yang et al., 2025b) utilizes scaled egocentric videos plus a unified human–robot action space with light robot finetuning, enabling efficient skill transfer and strong gains. Being-H0 (Luo et al., 2025) leverages physical-instruction tuning with discrete hand-motion codes (mm-level) and a physics-aligned cross-view space supports fine-grained VLA training from human videos. H-RDT (Bi et al., 2025) sets large bimanual pretraining with 3D hand pose and a two-stage, 2B-parameter diffusion policy delivers substantial improvements. GR-3 (Cheang et al., 2025) utilizes multi-source training (web, VR, robot trajectories) yields strong generalization, rapid few-shot adaptation, and robust long-horizon bimanual and mobile control. RynnVLA-001 (Jiang et al., 2025) pretrains on large-scale human video demonstrations with video generation objectives and compresses actions into a continuous latent space via ActionVAE to align video prediction with downstream robot fine-tuning. VITRA (Li et al., 2025c) treats the human hand as a proxy end-effector, converts in-the-wild egocentric hand videos into robot-aligned formats, and combines VLMs with diffusion-based action experts for policy learning.
These approaches rely on explicit alignment of human demonstrations to robot action spaces, which is inherently constrained by embodiment gaps between humans and robots. In contrast, our work targets a more upstream objective by transforming egocentric human data into embodiment supervision signals for an embodied brain, providing a scalable foundation that complements robot-data-based pipelines.
@@ 3.1 Egocentric2Embodiment Translation Pipeline
Our key idea is to translate egocentric human data into structured and verifiable supervision that captures the hierarchical structure of embodied behavior, spanning action semantics, temporal organization, interaction dynamics, and task-level reasoning. To this end, we design a schema-driven, rule-validated egocentric VQA data engine as shown in Fig.2 that systematically converts raw egocentric human videos into multi-level supervision aligned with embodied planning and interaction reasoning.
@@ 3.1.1 Data Intake and Pre-processing
To define the basic supervision units, the engine chunks each episode into short temporal clips, with episode-level metadata serving as contextual priors. Given the large variation in egocentric action amplitude and frequency across scenarios, we adopt scenario-aware temporal segmentation, including fixed-interval, event-driven, and kinematic-aware strategies. All clips are associated with explicit temporal spans and exposed through a unified interface for downstream annotation.
Episode-level metadata is used as contextual conditioning to limit the semantic space of subsequent question answering. The resulting representations are temporally localized and preserve short-range state transitions relevant to embodied manipulation and interaction.
3.1.2 Annotation Scheme Definition and Execution
To produce supervision that reflects embodied cognition rather than generic video description, we define a finite, schema-driven annotation space. Each clip is labeled with one of seven complementary VQA modes, including temporal, spatial, attribute, mechanics, reasoning, summary, and trajectory. Each mode is paired with a template set that standardizes wording and controls the information granularity. The engine samples a mode and a template, then generates a customized question and a detailed sentence answer for each clip.
VQA generation is performed by a set of VLM annotation engines. The schema constrains both the question form and the required semantic content, which keeps supervision targets consistent across different generators. Answers must be natural-language and grounded in the visual evidence. The engine enforces egocentric conventions such as left/right hand references and manipulation-specific phrasing such as contact verbs. This stage yields multi-level annotations that capture complementary aspects of planning and interaction reasoning.

### v2

@@ Abstract
Robotic generalization relies on physical intelligence: the ability to reason about state changes, contact-rich interactions, and long-horizon planning under egocentric perception and action. Vision Language Models (VLMs) are essential to Vision–Language–Action (VLA) systems, but the reliance on third-person training data creates a viewpoint gap for humanoid robots. Collecting massive robot-centric data is an ideal but impractical solution due to cost and diversity constraints. Conversely, human egocentric videos offer a highly scalable data source with rich interaction context, yet the embodiment mismatch prevents the direct application. To bridge this gap, we propose an Egocentric2Embodiment Translation Pipeline that transforms raw human egocentric videos into multi-level, schema-driven embodiment supervision with enforced evidence grounding and temporal consistency, enabling the construction of the Egocentric2Embodiment dataset (E2E-3M) at scale. An egocentric-aware embodied brain, termed PhysBrain, is obtained by training on the E2E-3M dataset. PhysBrain exhibits substantially improved egocentric understanding, particularly for planning. It provides an egocentric-aware initialization that enables more sample-efficient VLA fine-tuning and higher success rates, demonstrating effective transfer from human egocentric supervision to downstream robot control.
@@ 1 Introduction
Figure 1: Left: EgoThink radar plot comparing egocentric VLM performance across six dimensions. Right Top: The egocentric VLM performance on EgoThink and EgoPlan-B2 benchmarks with Qwen3VL-8B and PhysBrain, a Qwen3VL model fine-tuned with our annotated egocentric data (Sec. 3.1). Right Bottom: The egocentric VLA performance on SimplerEnv and RoboCasa benchmarks as the VLM backbone in a VLA fine-tuning pipeline.
@@
Vision-Language-Action (VLA) systems rely on a reliable embodied brain that integrates scenario understanding and action generation. Recent multimodal systems (Hurst et al., 2024a; Bai et al., 2025a) show rapid gains in visual perception, spatial and video reasoning, and long context understanding. These advances provide rich open vocabulary recognition and semantic inference capabilities that can be transferred to action prediction, thereby enabling modern VLAs (Zitkovich et al., 2023; Kim et al., 2024; Bjorck et al., 2025a; Black et al., 2024, 2025; Shen et al., 2025) to achieve strong performance across diverse manipulation tasks. These developments highlight that strong VLA performance is driven by an embodied brain that grounds executable planning and interaction decisions in the agent’s own perceptual stream.
@@
For future humanoid robots, this perceptual stream is expected to be predominantly first-person, since perception, planning, and action feasibility are fundamentally grounded in the agent’s own body and workspace (Grauman et al., 2022). This places stringent demands on multimodal models operating under egocentric settings. However, empirical results on egocentric benchmarks (Lin et al., 2022; Pramanick et al., 2023; Chen et al., 2024; Patel et al., 2025) indicate that current multimodal models still struggle with long-horizon understanding, planning, and reliability under egocentric videos as shown in Fig.1. These deficits stem from challenges intrinsic to egocentric perception, including rapid viewpoint changes, frequent hand–object occlusions, the absence of the actor’s full body, and the need for cross-frame inference of contact and object state (He et al., 2025). Consequently, current performance bottlenecks are more likely due to insufficient egocentric embodied cognition, state tracking, and planning supervision, rather than limitations in model scale or single-frame recognition.
@@
These limitations raise a fundamental scalability question: whether advancing VLA in egocentric settings necessarily depends on extensive robot data, including robot egocentric supervision. Acquiring large-scale and diverse robot manipulation data is widely acknowledged to be costly and difficult to scale, due to substantial hardware, labor, and safety constraints (Khazatsky et al., 2024). Even imitation learning relies on expensive human demonstrations, while existing large-scale robot data pipelines require long collection cycles or sustained multi-institution collaboration (Brohan et al., 2022; Zitkovich et al., 2023; O’Neill et al., 2024). As a result, learning and aligning embodied brains primarily through such robot data fundamentally constrains the scalability and coverage of egocentric VLA systems.
@@
In contrast to costly and hard-to-scale robot data, human first-person videos provide a naturally scalable source of egocentric supervision, covering diverse everyday behaviors and environments. This data modality offers observations closely aligned with real interaction distributions for learning embodied brains. Large-scale datasets, such as Ego4D (Grauman et al., 2022), BuildAI (BuildAI, 2025), and EgoDex (Hoque et al., 2025) demonstrate that egocentric videos can capture long-horizon activities, human–object interactions, and fine-grained manipulation dynamics at scale. An open question is how to leverage the latent planning structure and hand–object interaction regularities in human videos as supervision to strengthen egocentric embodied brains without robot data, thereby improving the sample efficiency and generalization of VLA systems.
@@
Motivated by this observation, we develop a scalable annotation and instruction pipeline: Egocentric2Embodiment (E2E) Translation Pipeline that transforms human egocentric videos into structured, multi-level first-person VQA supervision for embodied brain learning. Each VQA instance encodes complementary information across multiple levels, including planning decompositions, key states, interaction constraints, and temporal relations, providing supervision beyond static visual recognition.
We instantiate the E2E translation pipeline at scale to construct E2E-3M from Ego4D, BuildAI, and EgoDex. The resulting supervision is used to train PhysBrain over the Qwen3-VL backbone. As shown in Fig.1, PhysBrain substantially outperforms the base model on the Egocentric VLM benchmarks, with the clearest improvements in egocentric perception and planning. When plugged into the egocentric VLA system as the embodied brain, PhysBrain continues to get excellent performance on the VLA benchmarks. Few-shot adaptation on SimplerEnv and RoboCasa is effective and sample-efficient. The resulting performance exceeds VLA systems trained with large-scale robot data, while using no robot-data pretraining. Comprehensive evaluation results across VLM and VLA benchmarks validate E2E-3M dataset and the E2E Translation Pipeline, supporting human egocentric data as an effective source of physically grounded supervision for embodied brains. In summary, our contributions are as follows:
@@
  • We introduce a scalable annotation and instruction pipeline, called Egocentric2Embodiment Translation Pipeline, which converts large-scale human egocentric videos into multi-level embodied supervision.
@@ 2 Related Work
2.1 First-Person Vision Language Model
Vison Language Models (VLMs) that excel on third-person content often degrade when the input shifts to egocentric imagery and video. Multiple lines of evidence point to a persistent viewpoint domain gap and to missing egocentric cues such as hand manipulation, egomotion, and partial observability (He et al., 2025). EgoVLP (Lin et al., 2022) is among the first to document that third-person pretraining transfers poorly and that explicitly egocentric objectives are needed for first-person retrieval, recognition, and temporal grounding. EgoVLPv2 (Pramanick et al., 2023) further reports that fusing first-person video and language during pretraining is important for egocentric tasks. Beyond these early works, EgoPlan-Bench (Chen et al., 2024) shows that mainstream multimodal models struggle with egocentric planning even when the scenes are household and the instructions are simple, and it analyzes typical failure modes such as viewpoint confusion and missing contact reasoning. Studies on QaEgo4D (Barmann and Waibel, 2022) and QaEgo4Dv2 (Patel et al., 2025) find that both proprietary and open source VLMs lag on long-horizon egocentric reasoning. EgoM2P (Li et al., 2025a) also emphasizes the structural gap between third-person and first-person streams and argues for egocentric priors during pretraining.
@@ 2.2 Vision Language Action
Vision-Language-Action (VLA) models (Brohan et al., 2023; Zitkovich et al., 2023; Team et al., 2024; Black et al., 2024; Lian et al., 2026; Yu et al., 2026) represent a recent paradigm shift in robotic manipulation by unifying language understanding, visual perception, and motor control within a single end-to-end framework. Building upon large-scale vision-language models, VLAs directly map high-dimensional visual observations and natural language instructions to low-level robot actions, enabling intuitive human-robot interaction and task execution. Early works such as RT-1 (Brohan et al., 2023) and RT-2 (Zitkovich et al., 2023) demonstrate that scaling robot data and leveraging pretrained vision-language representations significantly improve manipulation performance across diverse tasks. Building upon these foundations, OpenVLA (Kim et al., 2024), $\pi_{0}$ (Black et al., 2024; Pertsch et al., 2025; Black et al., 2025), and GR00T-N1 (Bjorck et al., 2025a) further advance VLA capabilities through large-scale cross-embodiment and multi-task pretraining, demonstrating superior generalization and action prediction performance. Several works (Zhou et al., 2025; Yang et al., 2025d; Fang et al., 2025; Mazzaglia et al., 2025) attempt to address the catastrophic forgetting of language capabilities during VLA training, while others (Zawalski et al., 2025; Sun et al., 2024; Lin et al., 2025; Huang et al., 2025; Lee et al., 2025; Yuan et al., 2025) explore incorporating chain-of-thought reasoning into the VLA inference process. To pursue better generalization, several works (Shen et al., 2025; Liang et al., 2025; Jia et al., 2025) attempt to incorporate video generation models or world models into VLA action prediction, while others (Li et al., 2025b; Yu et al., 2025; Chen et al., 2025a, b) explore applying reinforcement learning to train VLA models.
@@ 2.3 Learning VLAs from Human Demonstration
Robot data acquisition is hard to scale due to the stringent robot–operator configuration and reliance on expert tele-operation. Egocentric VLA trained on the egocentric human demonstrations offers a more scalable path, with strong potential to advance perception–action learning and real-world executability. EgoVLA (Yang et al., 2025c) utilizes scaled egocentric videos plus a unified human–robot action space with light robot finetuning, enabling efficient skill transfer. Being-H0 (Luo et al., 2025) leverages physical-instruction tuning with discrete hand-motion codes and a physics-aligned cross-view space supports fine-grained VLA training from human videos. H-RDT (Bi et al., 2025) sets large bimanual pretraining with 3D hand pose and a two-stage diffusion policy delivers substantial improvements. GR-3 (Cheang et al., 2025) utilizes multi-source training yields strong generalization, rapid few-shot adaptation, and robust long-horizon bimanual and mobile control. RynnVLA-001 (Jiang et al., 2025) pretrains on large-scale human video demonstrations with video generation objectives and compresses actions into a continuous latent space via ActionVAE to align video prediction with downstream robot fine-tuning. VITRA (Li et al., 2025c) treats the human hand as a proxy end-effector, converts the egocentric hand videos into robot-aligned formats, and combines VLMs with diffusion-based action experts for policy learning.
However, these approaches focus on the explicit alignment of human data to robot action spaces, which is inherently constrained by embodiment gaps. In contrast, our work targets a more upstream objective by transforming egocentric human data into embodiment supervision signals for embodied brains, providing a scalable foundation that complements robot-data-based pipelines.
@@ 3.1 Egocentric2Embodiment Translation Pipeline
Our key idea is to translate human egocentric data into structured and verifiable supervision that captures the hierarchical structure of embodied behavior, spanning action semantics, temporal organization, interaction dynamics, and task-level reasoning. To this end, we design a schema-driven, rule-validated egocentric VQA data engine as shown in Fig.2 that systematically converts raw egocentric human videos into multi-level supervision aligned with embodied planning and interaction reasoning.
@@ 3.1.1 Data Intake and Pre-processing
To define the basic supervision units, the engine chunks each episode into short temporal clips at stage 1. Given the large variation in egocentric action amplitude and frequency across scenarios, we adopt scenario-aware temporal segmentation, including fixed-interval, event-driven(Build-AI, 2025), and kinematic-aware strategies(Li et al., 2025c). All clips are associated with explicit temporal spans and exposed through a unified interface for downstream annotation. Episode-level metadata is used as contextual conditioning to limit the semantic space of subsequent question answering.
The resulting representations are temporally localized and preserve short-range state transitions relevant to embodied manipulation and interaction. It forms the atomic unit for the downstream annotation stage.
3.1.2 Annotation Scheme
In stage 2, we define a finite, schema-driven annotation space to produce supervision that reflects embodied cognition rather than generic video description. Each clip is labeled with one of seven complementary VQA modes, including temporal, spatial, attribute, mechanics, reasoning, summary, and trajectory. Each mode is paired with a template set that standardizes wording and controls the information granularity. The engine samples a mode and a template, then generates a customized and detailed QA pair for each clip.
VQA generation is performed by a set of VLM annotation engines. The schema constrains both the question form and the required semantic content, which keeps supervision targets consistent across different generators. Answers must be natural-language and grounded in the visual evidence. The engine enforces egocentric conventions such as left/right hand references and manipulation-specific phrasing. This stage yields multi-level annotations that capture complementary aspects of planning and interaction reasoning.